mathslover (mathslover):

A train 100 metres long takes 6 seconds to cross a man walking at 5 kmph in the direction opposite to that of the train. Find the speed of the train.?

OpenStudy (kinggeorge):

So the man is going \(\frac{25}{18}\, m/s\). In 6 seconds, he moves exactly \[6\cdot\frac{25}{18}=\frac{25}{3} \,m\]Hence, the train goes \[100-\frac{25}{3}=\frac{275}{3}\,m\] in 6 seconds.

OpenStudy (kinggeorge):

That means the speed of the train should be\[\frac{275}{18}\,\,m/s\]

mathslover (mathslover):

yes in kmph it is 55

mathslover (mathslover):

55 kmph

mathslover (mathslover):

right sir ?

OpenStudy (kinggeorge):

Let me check.

mathslover (mathslover):

k sir

OpenStudy (kinggeorge):

Yup. It converts to 55 kph.

mathslover (mathslover):

so 55 is the correct answer right sir ?

OpenStudy (kinggeorge):

I believe so.

mathslover (mathslover):

thanks sir

OpenStudy (kinggeorge):

You're welcome.